# Covariance of Scattering Amplitudes from Counting Carefully

### v1

@@ 1 Introduction
Computing $\mathcal{A}_{a_{1}\dots a_{n}}$ for high values of $n$, which is a challenging task, has been tackled by several previous works, which this paper will draw upon. Combinatorial approaches have provided an iterative result for the ”tree sums”  Kreimer:2016jxo; Balduf:2019mai. However, the result does not distinguish the different possible index contractions among identical vertices. The path integral formalism has been used to derive a recursive method for the computation of Feynman rules Cohen:2023Covariance ; Cohen:2024Geometry ; cohen2025geometricbuildingblocks , which produce the amputated functions $\mathcal{A}_{a_{1}\dots a_{n}}$ up to ”evanescent” terms, i.e terms that vanish on-shell. The formalism has also been used to derive manifestly covariant Feynman rules Apostolos2020:Covariance ; HONERKAMP1972:Multiloops .
@@
  • An alternative combinatorial approach is adopted, based solely on the effective action $\Gamma[\phi]$ in combination with the vacuum and on-shell conditions. No assumptions are made regarding the type of fields involved nor the structure of the interaction terms, thus deriving the properties of the $\mathcal{S}$-matrix elements directly from the effective action, thus it holds for a generic EFT.
@@ 3.3 Counting topologies
Table 1: Number of Feynman diagrams W(n2)W(n-2) for $n=3,\dots,8$
Since Feynman diagrams are unrooted, the table counts using n2n-2 such that W(n2)W(n-2) corresponds to the number of Feynman diagrams with $n$ external legs.
@@
As mentioned earlier there are multiple combinations of vertices at the same level of gluing that contribute to the connected function such as in Fig.  3, which creates a need for going beyond the Ward numbers. A natural refinement can be obtained by examining a multivariate polynomial. The multivariate Ward polynomial is obtained by labeling each vertex by $a_{i-2}$ where $i$ is the number of legs HAIMAN1989:LagrangeInversion . For the 6 point case, corresponding to $W(4,k)$ it gives
@@
which can be seen as assigning weights to the the partitions identified in ( 11).
@@ 4 $a_{4}$ $10a_{2}^{2}+15a_{3}a_{1}$ $105a_{2}a_{1}^{2}$ $105a_{1}^{4}$
To recover the Ward polynomial simply evaluate $W_{n,k}(1,x,x^{2},\dots)$ and it is trivial to see that the right hand side fulfills the recursion relation defining the Ward numbers. The full tree sums given by $W_{n}$ can be obtained by evaluating $W_{n,k}(1,1,\dots)$
The multivariate Ward polynomial describes the Faà di Bruno FIGUEROA:HopfAlgebra ; Connes2000:Hopf1 and can be related to Feynman graphs  Connes2001:Renormalization ; Abdesselam2002:Feynman .
@@ 3.4 Refined Counting Function
$\sum_{k=1}^{n-2}\mathcal{F}_{n,k}(\vec{a}_{1},\vec{a}_{2},\dots\vec{a}_{n-k+1})x^{n-2}t^{2k-2}=a_{n,1}x^{n-2}$ (26)
$+\sum_{k=2}^{n-2}\frac{n!(k-2)!}{k!}\bigg(\prod_{j=1}^{k}\sum_{i=1}^{n-k-1}\sum_{l=1}^{i+2}\frac{a_{i+2,l}x^{i}}{(i+2-l)!(l-1)!}t^{l}\bigg)\bigg|_{x,t=0}$
@@ 4 On-Shell Connected Functions and Covariant Feynman Rules
The amputated connected function $\mathcal{A}_{n}$ is obtained by taking the sum over the different diagrams contributing to it, which are counted by ( 26).
@@
were introduced to simplify the expression. In addition to the transformation from the vertices, the propagators in (27) contribute a factor of $1/(\frac{\delta^{2}}{\delta\psi^{2}}\Gamma[\psi(\phi)]z_{1}^{2})^{k}$ when transforming. Additionally a factor of $(-1)^{k}$ is extracted, since each Feynman rule and propagator contributes a factor of $i$, which has been extracted explicitly.
@@
Sum over all transformation behaviors coming from gluing 1PI vertices to form $n$ point diagrams is then given by
@@
the same steps can be repeated for $f(x,y)$.
Evaluating the remaining derivatives and combining the terms leads to
@@ 4.1 Proof of covariance
combining (LABEL:First_sum_first_term) and (LABEL:second_sum) leads to
@@ 4.2 Existence of Covariant Feynman Rules
evaluating the derivatives with respect to $y$ then gives a factor of $\frac{(n+k-2)!}{(n+k-2-(k-2))!}=\frac{(n+k-2)!}{n!}$ resulting in

### v2

@@ 1 Introduction
Computing $\mathcal{A}_{a_{1}\dots a_{n}}$ for high values of $n$, which is a challenging task, has been tackled by several previous works, which this paper will draw upon. Combinatorial approaches have provided an iterative result for the ”tree sums”  Kreimer:Combinatorics2016 ; Balduf:Combinatorics2020 . However, the result does not distinguish the different possible index contractions among identical vertices. The path integral formalism has been used to derive a recursive method for the computation of Feynman rules Cohen:2023Covariance ; Cohen:2024Geometry ; cohen2025geometricbuildingblocks , which produce the amputated functions $\mathcal{A}_{a_{1}\dots a_{n}}$ up to ”evanescent” terms, i.e terms that vanish on-shell. The formalism has also been used to derive manifestly covariant Feynman rules Apostolos2020:Covariance ; HONERKAMP1972:Multiloops .
@@
  • An alternative combinatorial approach is adopted, based solely on the effective action $\Gamma[\phi]$ in combination with the vacuum and on-shell conditions. No assumptions are made regarding the type of fields involved nor the structure of the interaction terms, thus deriving the properties of the $\mathcal{S}$-matrix elements directly from the effective action, hence it holds for a generic EFT.
@@ 3.3 Counting topologies
Table 1: Number of Feynman diagrams W(n2)W(n-2) for $n=2,\dots,7$
Since Feynman diagrams are unrooted, the table counts using n2n-2 such that W(n2)W(n-2) corresponds to the number of Feynman diagrams with $n$ external legs, starting from $n=2$ as the porpagator.
@@
As mentioned earlier there are multiple combinations of vertices at the same level of gluing that contribute to the connected function such as in Fig. 3, which creates a need for going beyond the Ward numbers. A natural refinement can be obtained by examining a multivariate polynomial. The multivariate Ward polynomial is obtained by labeling each vertex by $a_{i-2}$ where $i$ is the number of legs HAIMAN1989:LagrangeInversion . For the 6 point case, corresponding to $W(4,k)$ it gives
@@
which can be seen as assigning weights to the the partitions identified in (11).
@@ 4 $a_{4}$ $10a_{2}^{2}+15a_{3}a_{1}$ $105a_{2}a_{1}^{2}$ $105a_{1}^{4}$
To recover the Ward polynomial simply evaluate $W_{n,k}(1,x,x^{2},\dots)$ and it is trivial to see that the right hand side fulfills the recursion relation defining the Ward numbers. The full tree sums given by $W_{n}$ can be obtained by evaluating $W_{n,k}(1,1,\dots)$. The multivariate Ward polynomial describes the Faà di Bruno algebra FIGUEROA:HopfAlgebra ; Connes2000:Hopf1 and can be related to Feynman graphs  Connes2001:Renormalization ; Abdesselam2002:Feynman .
@@ 3.4 Refined Counting Function
$\sum_{k=1}^{n-2}\mathcal{F}_{n,k}(\vec{a}_{1},\vec{a}_{2},\dots\vec{a}_{n-k+1})x^{n-2}t^{2k-2}=$ (26)
$a_{n,0}x^{n-2}+\sum_{k=2}^{n-2}\frac{n!(k-2)!}{k!}\bigg(\prod_{j=1}^{k}\sum_{i=1}^{n-k-1}\sum_{l=1}^{i+2}\frac{a_{i+2,l}x^{i}}{(i+2-l)!(l-1)!}t^{l}\bigg)$
@@ 4 On-Shell Connected Functions and Covariant Feynman Rules
The amputated connected function $\mathcal{A}_{n}$ is obtained by taking the sum over the different diagrams contributing to it, which are counted by (26).
@@
were introduced to simplify the expression. In addition to the transformation from the vertices, the propagators in (27) contribute a factor of $1/(\frac{\delta^{2}}{\delta\psi^{2}}\Gamma[\psi(\phi)]z_{1}^{2})^{k-1}$ when transforming. Additionally a factor of $(-1)^{k-1}$ emerges, since each Feynman rule and propagator contributes a factor of $i$, which has been extracted explicitly.
@@
The sum over all transformation behaviors coming from gluing 1PI vertices to form $n$ point diagrams is then given by
@@
the same steps can be repeated for $f(x,y)$. Evaluating the remaining derivatives and combining the terms leads to
@@ 4.1 Proof of covariance
combining (59) and (LABEL:second_sum) leads to
@@ 4.2 Existence of Covariant Feynman Rules
evaluating the derivatives with respect to $y$ then gives a factor of (n+k2)!/(n+k2(k2))!=(n+k2)!/n!{(n+k-2)!}/{(n+k-2-(k-2))!}={(n+k-2)!}/{n!} resulting in